OpenStudy (anonymous):

Verify the identity cot[ theta - pi/2 ] = -tan theta

ganeshie8 (ganeshie8):

take left hand side expression, change it into sin and cos

ganeshie8 (ganeshie8):

use below : cot = cos/sin

OpenStudy (anonymous):

\[\cot \left[ \theta - \frac{ \pi }{ 2 } \right] = -\tan \theta\] Where do I get cos and sin from o.o

ganeshie8 (ganeshie8):

^^ second reply

ganeshie8 (ganeshie8):

Left hand side : \(\large \cot \left[ \theta - \frac{ \pi }{ 2 } \right]\) \(\large \frac{\cos \left[ \theta - \frac{ \pi }{ 2 } \right]}{\sin \left[ \theta - \frac{ \pi }{ 2 } \right]}\)

OpenStudy (anonymous):

So whenever I want to convert cot to cos and sin I just add that to the numerator and denominator?

ganeshie8 (ganeshie8):

by definition, \(\large \cot x = \frac{\cos x}{\sin x}\)

ganeshie8 (ganeshie8):

and yes, where ever u see "cot", u can replace it wid "cos/sin"

ganeshie8 (ganeshie8):

next use below identities :- cos(-x) = cos(x) sin(-x) = -sin(x) cos(pi/2-x) = sin(x) sin(pi/2-x) = cos(x)

ganeshie8 (ganeshie8):

Left hand side : \(\large \cot \left[ \theta - \frac{ \pi }{ 2 } \right]\) \(\large \frac{\cos \left[ \theta - \frac{ \pi }{ 2 } \right]}{\sin \left[ \theta - \frac{ \pi }{ 2 } \right]} \) \(\large \frac{\cos \left[ -(\frac{ \pi }{ 2 } -\theta) \right]}{\sin \left[- (\frac{ \pi }{ 2 }-\theta) \right]} \) \(\large \frac{\cos \left[ \frac{ \pi }{ 2 } -\theta \right]}{-\sin \left[ (\frac{ \pi }{ 2 }-\theta) \right]} \)

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yesh...

ganeshie8 (ganeshie8):

good :) next use the identites : cos(pi/2-x) = sin(x) sin(pi/2-x) = cos(x)

ganeshie8 (ganeshie8):

Left hand side : \(\large \cot \left[ \theta - \frac{ \pi }{ 2 } \right]\) \(\large \frac{\cos \left[ \theta - \frac{ \pi }{ 2 } \right]}{\sin \left[ \theta - \frac{ \pi }{ 2 } \right]} \) \(\large \frac{\cos \left[ -(\frac{ \pi }{ 2 } -\theta) \right]}{\sin \left[- (\frac{ \pi }{ 2 }-\theta) \right]} \) \(\large \frac{\cos \left[ \frac{ \pi }{ 2 } -\theta \right]}{-\sin \left[ (\frac{ \pi }{ 2 }-\theta) \right]} \) \(\large \frac{\sin \theta}{-\cos \theta} \) \(\large -\tan \theta \) QED

ganeshie8 (ganeshie8):

let me knw if domething doesnt make sense..

OpenStudy (anonymous):

How come theta doesn't get canceled out

OpenStudy (anonymous):